Ref # 1102550
----------
cos(x) = 2
Solve for x.
-------------
As an example from my book:
cos(+/-arctan(2sqrt(6)/5) - i*ln(7) + n*2pi) = 10
=======================================================
--> cos(x) = 10
x = +/-arctan(2sqrt(6)/5) - i*ln(7) + n*2pi
============================
If cos(x) = 10 has a solution,
I think cos(x) = 2 does also.
I solved that years ago, but I don't remember how to do it now.
I will figure it out, with or without help.
================
Then I wonder about cos(x) = 2 + i ???

Answer by rothauserc(4718)

cos(x) = 2 + i
:
take the inverse cosine of both sides to eliminate the cosine from the left hand side.
:
********************************************************
x = 2 * π * n(1) + cos^(-1)(2 + i) for n(1) element of Z
:
or x = 2 * π * n(2) - cos^(-1)(2 + i) for n(2) element of Z
:
Z is the set of integers
********************************************************
: